OpenStudy (anonymous):

Let f(x)= sqrt(sin(3x)). Find f'(x)

OpenStudy (anonymous):

chain rule it all the way down

OpenStudy (anonymous):

Idk the derivative of sqrt. Lead me through?

OpenStudy (anonymous):

\[\frac{d}{dx}[\sqrt{f(x)}]=\frac{f'(x)}{2\sqrt{f(x)}}\] with \[f(x)=\sin(3x), f'(x)=3\cos(3x)\]

OpenStudy (anonymous):

whoah hold the phone you don't know the derivative of the square root function?

OpenStudy (anonymous):

no. lol

OpenStudy (anonymous):

it is a very very common function you should just memorize it, kinda like knowing \(8\times 9=72\) if \[f(x)=\sqrt{x}\] then \[f'(x)=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

it never changes so no need to do something ridiculous like writing \[\sqrt{x}=x^{\frac{1}{2}}\] \[\frac{d}{dx}[x^{\frac{1}{2}}]=\frac{1}{2}x^{\frac{1}{2}-1}=\frac{1}{2}x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

just know that the derivative of the square root of x is one over two root x that is all

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and appying that with the chain rule tells you that \[\frac{d}{dx}\sqrt{g(x)}=\frac{g'(x)}{2\sqrt{g(x)}}\]

OpenStudy (anonymous):

So I got \[1/2\sqrt{x} (\sin(3x))+xos(3x)+3\]...?

OpenStudy (anonymous):

on no dear not plus, times

OpenStudy (anonymous):

this is all wrong, lets go slow

OpenStudy (anonymous):

Ohhh. Sorry. I kinda missed a day of school when we were taught this.

OpenStudy (anonymous):

ok in english the derivative of the square root of x is one over two root x the derivative of the square root of g of x is one over two root g of x, times the derivative of g

OpenStudy (anonymous):

in math \[\frac{d}{dx}\sqrt{g(x)}]=\frac{g'(x)}{2\sqrt{g(x)}}\]

OpenStudy (anonymous):

in your case you have \[f(x)=\sqrt{\sin(3x)}\]right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so the first step is to write one over two root sine of 3 x

OpenStudy (anonymous):

Why is it one?

OpenStudy (anonymous):

i.e. \[\frac{1}{2\sqrt{\sin(3x)}}\times something \]

OpenStudy (anonymous):

it will not be one when we are done

OpenStudy (anonymous):

what goes on top is the derivative of \(\sin(3x)\)

OpenStudy (anonymous):

So that's temporary for now?

OpenStudy (anonymous):

zactly now what is the derivative of \(\sin(3x)\)?

OpenStudy (anonymous):

cos(3x)*3?

OpenStudy (anonymous):

yes, which most would write as \[3\cos(3x)\] putting the coefficient out front that is what goes in the numerator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

final answer? \[\frac{3\cos(3x)}{2\sqrt{\sin(3x)}}\] done

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

ok not really yes, that is all

OpenStudy (anonymous):

oh okay. haha. Thank y ou!

OpenStudy (anonymous):

yw